Janet sells flowers that cost her $2.25 each. She assumes that 10% of these will go bad and have to be thrown away. Janet wants a 35% markup based on cost and buys 200 of these flowers. What should Janet charge Jack for each flower he buys?
Answer by ankor@dixie-net.com(22740)

Janet sells flowers that cost her $2.25 each.
She assumes that 10% of these will go bad and have to be thrown away.
Janet wants a 35% markup based on cost and buys 200 of these flowers.
What should Janet charge Jack for each flower he buys?
:
Find the cost of the 200 flowers
2.25 * 200 = $450
A 35% mark-up: 1.35 * 450 = $607.50 revenue desired
:
It says 10% go bad, 90% are OK: .9(200) = 180 flowers to sell
Derive this amt from selling 180 flowers
607.50/180 = $3.375 ~ $3.38 charged for each flower